# A new (but very nearly old) proof of the Rogers–Ramanujan identities

### v2

@@
Abstract.
@@
Supported by the Swedish Science Research Council, project no. 2020-04221.
1. Introduction
@@
$\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{(1-q)(1-q^{2})\dotsm(1-q^{k})}$ $=\prod_{j=0}^{\infty}\frac{1}{(1-q^{5j+1})(1-q^{5j+4})},$ (1.1a)
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(1-q)(1-q^{2})\dotsm(1-q^{k})}$ $=\prod_{j=0}^{\infty}\frac{1}{(1-q^{5j+2})(1-q^{5j+3})}$ (1.1b)
has been told many times. They were first proved by Rogers in a brilliant but forgotten paper from 1894 and rediscovered by Ramanujan some time before 1913. Hardy wrote about the seven proofs known in the 1930s: “None of them is both simple and straight-forward, and probably it would be unreasonable to hope for a proof which is” [H]. Several dozen more proofs have appeared since then, using diverse methods. We refer to [A2] for a survey of the proofs available in the 1980s and [AKP, BP, C, F, GIS, P, St, T] for examples of more recent work. The purpose of the present note is to give a new proof that, although far from the shortest possible, we find relatively “simple and straight-forward”. Almost unbelievably, this proof was very nearly found by Rogers. He knew all its main ingredients and could easily have given it by adding one or two sentences to [R2].
Among the many gems buried in Rogers’ paper are the identities [R2, §6, (5)–(8)]
$G(q)+G(-q)$ $=2F(q)G(q^{16}),$ (1.2a)
$G(q)-G(-q)$ $=2qF(q)H(-q^{4}),$ (1.2b)
$H(q)+H(-q)$ $=2F(q)G(-q^{4}),$ (1.2c)
$H(q)-H(-q)$ $=2q^{3}F(q)H(q^{16}),$ (1.2d)
@@
Below, we will give both a version of Rogers’ proof of (1) and another proof based on work of Andrews [A1]. After deriving (1), Rogers writes somewhat dismissively
@@
At this point Rogers has already proved the Rogers–Ramanujan identities. He realizes that (1) are quite easy to prove for the product sides of (1) and hence perhaps less interesting than they may seem.
Our observation is simply that the identities (1) uniquely determine $G$ and $H$ up to normalization. In fact, if they are written as
$G(q)$ $=F(q)\left(G(q^{16})+qH(-q^{4})\right),$ (1.3a)
$H(q)$ $=F(q)\left(G(-q^{4})+q^{3}H(q^{16})\right),$ (1.3b)
they can be used to recursively compute $G$ and $H$ as power series, starting from the initial values $G(0)=H(0)=1$. Since Rogers has proved that (1) hold both for the sum sides and the product sides of (1), this gives an independent proof of the Rogers–Ramanujan identities, which is considerably easier than the one given earlier in the same paper.
We will present the resulting proof in some detail, but we stress that apart from the preceding paragraph everything can be found in [A1] and [R2]. I am very much indebted to George Andrews for pointing this out to me. I would also like to thank Andrew Sills for the reference to his master’s thesis and for making it available [S1].
In §3, we explain how some further known Rogers–Ramanujan-type identities either appear as byproducts of the proof or can be obtained by slight variations. We only consider identities that are very close to (1), in that the product sides are essentially the same. It would be interesting to investigate whether similar proofs can be given for other results. As a first step, in [Ro] we show that the method can be used to prove Rogers’ and Slater’s mod $7$ identities.
2. A proof of the Rogers–Ramanujan identities
@@
$(a;q)_{k}=\prod_{j=0}^{k-1}(1-aq^{j}),\qquad k\in\mathbb{Z}_{\geq 0}\cup\{\infty\},$
which is extended to negative indices by $(a;q)_{-k}=1/(aq^{-k};q)_{k}$. More generally,
@@
and
$G(q)=R(1;q),\qquad H(q)=R(q;q).$
The Rogers–Ramanujan identities (1) then take the compact form
$G(q)=\frac{1}{(q,q^{4};q^{5})_{\infty}},\qquad H(q)=\frac{1}{(q^{2},q^{3};q^{5})_{\infty}}.$ (2.1)
Finally, we introduce the multiplicative theta function
$\theta(z;q)=(q,z,q/z;q)_{\infty}.$
@@
To some readers these identities may look just as complicated as (1), but they are much easier to prove. For instance, they can all be obtained as special cases of Ramanujan’s ${}_{1}\psi_{1}$-summation, a simple proof of which was given by Ismail [I], see also [A1, App. C].
Following Andrews [A1], we will represent the sums in (1) as constant terms. In general, if
$f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}$
(all series in this paper can be considered as formal or analytic according to taste) and we let $\operatorname{CT}$ denote the constant term of a Laurent series in $z$, then
$\operatorname{CT}\theta(1/z;Q)f(xz^{N})$ $=\operatorname{CT}\sum_{l=-\infty}^{\infty}(-1)^{l}Q^{l(l-1)/2}z^{-l}\sum_{k=0}^{\infty}a_{k}x^{k}z^{Nk}$
@@
$R(t;q)=\operatorname{CT}\frac{\theta(1/z;q^{2})}{(-qtz;q)_{\infty}}=\operatorname{CT}\frac{\theta(1/z;q^{1/2})}{(q^{1/2}tz^{2};q)_{\infty}}.$
@@
$\frac{\theta(1/z;q^{2})}{(-qtz;q)_{\infty}}$ $=\frac{\theta(1/z;q^{2})}{(-qtz;q^{2})}\cdot\frac{1}{(-q^{2}tz;q^{2})}=\frac{\theta(1/z;q^{2})}{(-q^{2}tz;q^{2})}\cdot\frac{1}{(-qtz;q^{2})},$
$\frac{\theta(1/z;q^{1/2})}{(q^{1/2}tz^{2};q)_{\infty}}$ $=\frac{\theta(1/z;q^{1/2})}{(q^{1/4}t^{1/2}z;q^{1/2})_{\infty}}\cdot\frac{1}{(-q^{1/4}t^{1/2}z;q^{1/2})_{\infty}},$
@@
$R(t;q)$ $=(-qt;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}t^{k}}{(q^{2},-qt;q^{2})_{k}}$ (2.6a)
$=(-q^{2}t;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}}t^{k}}{(q^{2},-q^{2}t;q^{2})_{k}}$ (2.6b)
$=(q^{1/4}t^{1/2};q^{1/2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}/4}t^{k/2}}{(q^{1/2},q^{1/4}t^{1/2};q^{1/2})_{k}}.$ (2.6c)
These identities were found by Rogers [R2, §6, (2)–(4)]. Except for his way of writing, the original proof is not difficult; we will review it in §4. Watson independently proved (2.6b) and (2.6c) [W].
For $t=1$ and $t=q$, the equations (2) simplify to
$G(q)$ $=(-q;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(-q;-q)_{2k}}$ (2.7a)
$=(-q^{2};q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{(q^{4};q^{4})_{k}}$ (2.7b)
$=(q^{1/4};q^{1/2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{{k^{2}}/4}}{(q^{1/4};q^{1/4})_{2k}},$ (2.7c)
$H(q)$ $=(-q^{2};q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+2)}}{(q^{4};q^{4})_{k}}$ (2.7d)
$=(-q;q^{2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(-q;-q)_{2k+1}}$ (2.7e)
$=(q^{1/4};q^{1/2})_{\infty}\sum_{k=0}^{\infty}\frac{q^{{k(k+2)}/4}}{(q^{1/4};q^{1/4})_{2k+1}}.$ (2.7f)
@@
$A(q)=\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{(q^{4};q^{4})_{k}},\qquad B(q)=\sum_{k=0}^{\infty}\frac{q^{k(k+2)}}{(q^{4};q^{4})_{k}}.$ (2.8)
Then, (2) can be written
$G(-q^{4})$ $=(q^{4};q^{8})_{\infty}\frac{B(q)+B(-q)}{2},$ (2.9a)
$G(q)$ $=(-q^{2};q^{2})_{\infty}\,A(q),$ (2.9b)
$G(q^{16})$ $=(q^{4};q^{8})_{\infty}\frac{A(q)+A(-q)}{2},$ (2.9c)
$H(q)$ $=(-q^{2};q^{2})_{\infty}\,B(q),$ (2.9d)
$H(-q^{4})$ $=(q^{4};q^{8})_{\infty}\frac{A(q)-A(-q)}{2q},$ (2.9e)
$H(q^{16})$ $=(q^{4};q^{8})_{\infty}\frac{B(q)-B(-q)}{2q^{3}}.$ (2.9f)
We may use (2.9b) and (2.9d) to eliminate $A$ and $B$ from the remaining equations. Writing also
$F(q)=\frac{(-q^{2};q^{2})_{\infty}}{(q^{4};q^{8})_{\infty}}=\frac{(q^{8};q^{8})_{\infty}}{(q^{2};q^{2})_{\infty}}$
we arrive at (1) or, equivalently, (1).
As was noted in the introduction, the identities (1) can be considered as recursions for computing the Taylor series of $G$ and $H$, starting from $G(0)=H(0)=1$; see (2) below. Hence, if we can verify (1) for the product sides
$G(q)=\frac{1}{(q,q^{4};q^{5})_{\infty}},\qquad H(q)=\frac{1}{(q^{2},q^{3};q^{5})_{\infty}},$
the Rogers–Ramanujan identities follow. As Rogers observed, this reduces to classical theta function identities. We first write
$G(q)$ $=\frac{1}{(q,q^{4},q^{6},q^{9};q^{10})_{\infty}}=\frac{(-q,-q^{9};q^{10})_{\infty}}{(q^{4},q^{6};q^{10})_{\infty}(q^{2},q^{18};q^{20})_{\infty}}$
$=\frac{(q^{8},q^{12};q^{20})_{\infty}\,\theta(-q;q^{10})}{(q^{2};q^{2})_{\infty}}$
@@
$H(q)=\frac{(q^{4},q^{16};q^{20})_{\infty}\,\theta(-q^{3};q^{10})}{(q^{2};q^{2})_{\infty}}.$
@@
$\frac{\theta(z;q)+\theta(-z;q)}{2}$ $=\sum_{k=-\infty}^{\infty}q^{k(2k-1)}z^{2k}=\theta(-qz^{2};q^{4}),$
$\frac{\theta(z;q)-\theta(-z;q)}{2}$ $=-\sum_{k=-\infty}^{\infty}q^{k(2k+1)}z^{2k+1}=-z\,\theta(-q^{3}z^{2};q^{4}).$
@@
$\frac{G(q)+G(-q)}{2}=\frac{(q^{8},q^{12};q^{20})_{\infty}(-q^{12},-q^{28},q^{40};q^{40})_{\infty}}{(q^{2};q^{2})_{\infty}}.$
@@
$(-q^{12},-q^{28};q^{40})_{\infty}=\frac{(q^{24},q^{56};q^{80})_{\infty}}{(q^{12},q^{28};q^{40})_{\infty}}=\frac{(q^{16},q^{24};q^{40})_{\infty}}{(q^{12},q^{28};q^{40})_{\infty}(q^{16},q^{64};q^{80})}$
@@
$\frac{G(q)+G(-q)}{2}=\frac{(q^{8};q^{8})_{\infty}}{(q^{2};q^{2})_{\infty}(q^{16},q^{64};q^{80})_{\infty}}=\frac{(q^{8};q^{8})_{\infty}}{(q^{2};q^{2})_{\infty}}\,G(q^{16}).$
The remaining parts of (1) follow similarly. Alternatively, the identities (1) can be obtained directly as special cases of
$\theta(z;q)=\theta(-qz^{2};q^{4})-z\,\theta(-q^{3}z^{2};q^{4}).$
@@
$F(q)=\sum_{k=0}^{\infty}f_{k}q^{2k},\quad G(q)=\sum_{k=0}^{\infty}g_{k}q^{k},\quad H(q)=\sum_{k=0}^{\infty}h_{k}q^{k}$
into (1) gives
$g_{2k}$ $=\sum_{j=0}^{\lfloor k/8\rfloor}f_{k-8j}g_{j},$ $g_{2k+1}$ $=\sum_{j=0}^{\lfloor k/2\rfloor}(-1)^{j}f_{k-2j}h_{j},$ (2.10a)
$h_{2k}$ $=\sum_{j=0}^{\lfloor k/2\rfloor}(-1)^{j}f_{k-2j}g_{j},$ $h_{2k+1}$ $=\sum_{j=0}^{\lfloor(k-1)/8\rfloor}f_{k-8j-1}g_{j}.$ (2.10b)
It is interesting to note that all the coefficients involved here have partition-theoretic interpretations. This is straight-forward for $f_{n}$, and very well-known for $g_{n}$ and $h_{n}$. This leads to the following result, which can be viewed as a combinatorial reformulation of (1).
@@
Let $f_{n}$ be the number of partitions of $n$ into parts not divisible by $4$. Let $g_{n}$ be the number of partitions of $n$ into parts congruent to ±$\pm 1$ $\operatorname{mod}$ $5$ or, equivalently, the number of partitions of $n$ without repeated or consecutive parts. Let $h_{n}$ be the number of partitions of $n$ into parts congruent to ±$\pm 2$ $\operatorname{mod}$ $5$ or, equivalently, the number of partitions of $n$ without repeated or consecutive parts and not containing $1$ as a part. Then, the recursions (2) hold.
3. Some additional results
Having proved the Rogers–Ramanujan identities, we can sum the six series in (2). This was of course clear to Rogers, and leads to the identities given in Slater’s list [MSZ, S2, Sl] as number 99, 20, 98, 16, 94 and 96, respectively (as is noted in [S2], 98 is equivalent to 79, and 94 to 17).

### v3

@@
[…]
A New (But Very Nearly Old) Proof of the Rogers–Ramanujan Identities
[…]
A New (But Very Nearly Old) Proof
of the Rogers–Ramanujan IdentitiesThis paper is a contribution to the Special Issue on Basic Hypergeometric Series Associated with Root Systems and Applications in honor of Stephen C. Milne’s 75th birthday. The full collection is available at https://www.emis.de/journals/SIGMA/Milne.html
[…]
Hjalmar ROSENGREN
[…]
H. Rosengren
[…]
Department of Mathematical Sciences, Chalmers University of Technology
[…]
Received March 25, 2024, in final form June 27, 2024; Published online July 02, 2024
[…]
@@ H. Rosengren
[…]
Rogers–Ramanujan identities; constant term identities
[…]
11P84; 33D15
Dedicated to Stephen Milne on his 75th birthday
1 Introduction
@@
$\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{(1-q)\bigl{(}1-q^{2}\bigr{)}\dotsm(1-q^{k})}$ $=\prod_{j=0}^{\infty}\frac{1}{\bigl{(}1-q^{5j+1}\bigr{)}\bigl{(}1-q^{5j+4}\bigr{)}},$ (1.1a)
$\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(1-q)(1-q^{2})\dotsm\bigl{(}1-q^{k}\bigr{)}}$ $=\prod_{j=0}^{\infty}\frac{1}{\bigl{(}1-q^{5j+2}\bigr{)}\bigl{(}1-q^{5j+3}\bigr{)}}$ (1.1b)
has been told many times. They were first proved by Rogers in a brilliant but forgotten paper from 1894 and rediscovered by Ramanujan some time before 1913. Hardy wrote about the seven proofs known in the 1930s: “None of them is both simple and straight-forward, and probably it would be unreasonable to hope for a proof which is” [10]. Several dozen more proofs have appeared since then, using diverse methods. We refer to [2] for a survey of the proofs available in the 1980s and [3, 5, 6, 7, 8, 13, 21, 22] for examples of more recent work. The purpose of the present note is to give a new proof that, although far from the shortest possible, we find relatively “simple and straight-forward”. Almost unbelievably, this proof was very nearly found by Rogers. He knew all its main ingredients and could easily have given it by adding one or two sentences to [15]. Among the many gems buried in Rogers’ paper are the identities [15, Section 6, formulas (5)–(8)]
$G(q)+G(-q)=2F(q)G\bigl{(}q^{16}\bigr{)},\qquad$ $G(q)-G(-q)=2qF(q)H\bigl{(}-q^{4}\bigr{)},$ (1.2a)
$H(q)+H(-q)=2F(q)G\bigl{(}-q^{4}\bigr{)},\qquad$ $H(q)-H(-q)=2q^{3}F(q)H\bigl{(}q^{16}\bigr{)},$ (1.2b)
@@
Below, we will give both a version of Rogers’ proof of (1.2) and another proof based on work of Andrews [1]. After deriving (1.2), Rogers writes somewhat dismissively
@@
At this point Rogers has already proved the Rogers–Ramanujan identities. He realizes that (1.2) are quite easy to prove for the product sides of (1.1) and hence perhaps less interesting than they may seem.
Our observation is simply that the identities (1.2) uniquely determine $G$ and $H$ up to normalization. In fact, if they are written as
$G(q)=F(q)\bigl{(}G\bigl{(}q^{16}\bigr{)}+qH\bigl{(}-q^{4}\bigr{)}\bigr{)},$ (1.3a)
$H(q)=F(q)\bigl{(}G\bigl{(}-q^{4}\bigr{)}+q^{3}H\bigl{(}q^{16}\bigr{)}\bigr{)},$ (1.3b)
they can be used to recursively compute $G$ and $H$ as power series, starting from the initial values $G(0)=H(0)=1$. Since Rogers has proved that (1.2) hold both for the sum sides and the product sides of (1.1), this gives an independent proof of the Rogers–Ramanujan identities, which is considerably easier than the one given earlier in the same paper.
We will present the resulting proof in some detail, but we stress that apart from the preceding paragraph everything can be found in [1, 15]. I am very much indebted to George Andrews for pointing this out to me.
In Section 3, we explain how some further known Rogers–Ramanujan-type identities either appear as byproducts of the proof or can be obtained by slight variations. We only consider identities that are very close to (1.1), in that the product sides are essentially the same. It would be interesting to investigate whether similar proofs can be given for other results. As a first step, in [17] we show that the method can be used to prove Rogers’ and Slater’s mod $7$ identities.
2 A proof of the Rogers–Ramanujan identities
@@
$(a;q)_{k}=\prod_{j=0}^{k-1}\bigl{(}1-aq^{j}\bigr{)},\qquad k\in\mathbb{Z}_{\geq 0}\cup\{\infty\},$
which is extended to negative indices by $(a;q)_{-k}=1/\bigl{(}aq^{-k};q\bigr{)}_{k}$. More generally,
@@
and $G(q)=R(1;q)$, $H(q)=R(q;q)$. The Rogers–Ramanujan identities (1.1) then take the compact form
$G(q)=\frac{1}{\bigl{(}q,q^{4};q^{5}\bigr{)}_{\infty}},\qquad H(q)=\frac{1}{\bigl{(}q^{2},q^{3};q^{5}\bigr{)}_{\infty}}.$
Finally, we introduce the multiplicative theta function $\theta(z;q)=(q,z,q/z;q)_{\infty}$.
@@
To some readers these identities may look just as complicated as (1.1), but they are much easier to prove. For instance, they can all be obtained as special cases of Ramanujan’s ${}_{1}\psi_{1}$-summation, a simple proof of which was given by Ismail [11], see also [1, Appendix C].
Following Andrews [1], we will represent the sums in (1.1) as constant terms. In general, if ${f(z)=\sum_{k=0}^{\infty}a_{k}z^{k}}$ (all series in this paper can be considered as formal or analytic according to taste) and we let $\operatorname{CT}$ denote the constant term of a Laurent series in $z$, then by (2.2), we have
$\operatorname{CT}\theta(1/z;Q)f\bigl{(}xz^{N}\bigr{)}$ $=\operatorname{CT}\sum_{l=-\infty}^{\infty}(-1)^{l}Q^{l(l-1)/2}z^{-l}\sum_{k=0}^{\infty}a_{k}x^{k}z^{Nk}$
@@
$R(t;q)=\operatorname{CT}\frac{\theta\bigl{(}1/z;q^{2}\bigr{)}}{(-qtz;q)_{\infty}}=\operatorname{CT}\frac{\theta\bigl{(}1/z;q^{1/2}\bigr{)}}{\bigl{(}q^{1/2}tz^{2};q\bigr{)}_{\infty}}.$
@@
$\frac{\theta\bigl{(}1/z;q^{2}\bigr{)}}{(-qtz;q)_{\infty}}=\frac{\theta\bigl{(}1/z;q^{2}\bigr{)}}{\bigl{(}-qtz;q^{2}\bigr{)}_{\infty}}\cdot\frac{1}{\bigl{(}-q^{2}tz;q^{2}\bigr{)}_{\infty}}=\frac{\theta(1/z;q^{2})_{\infty}}{\bigl{(}-q^{2}tz;q^{2}\bigr{)}_{\infty}}\cdot\frac{1}{\bigl{(}-qtz;q^{2}\bigr{)}_{\infty}},$
$\frac{\theta\bigl{(}1/z;q^{1/2}\bigr{)}}{\bigl{(}q^{1/2}tz^{2};q\bigr{)}_{\infty}}=\frac{\theta\bigl{(}1/z;q^{1/2}\bigr{)}}{\bigl{(}q^{1/4}t^{1/2}z;q^{1/2}\bigr{)}_{\infty}}\cdot\frac{1}{\bigl{(}-q^{1/4}t^{1/2}z;q^{1/2}\bigr{)}_{\infty}},$
@@
$R(t;q)$ $=\bigl{(}-qt;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}t^{k}}{\bigl{(}q^{2},-qt;q^{2}\bigr{)}_{k}}$ (2.5a)
$=\bigl{(}-q^{2}t;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}}t^{k}}{\bigl{(}q^{2},-q^{2}t;q^{2}\bigr{)}_{k}}$ (2.5b)
$=\bigl{(}q^{1/4}t^{1/2};q^{1/2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}/4}t^{k/2}}{\bigl{(}q^{1/2},q^{1/4}t^{1/2};q^{1/2}\bigr{)}_{k}}.$ (2.5c)
These identities were found by Rogers [15, Section 6, formulas (2)–(4)]. Except for his way of writing, the original proof is not difficult; we will review it in Section 4. Watson independently proved (2.5b) and (2.5c) [23].
For $t=1$ and $t=q$, the equations (2.5) simplify to
$G(q)$ $=\bigl{(}-q;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(-q;-q)_{2k}}$ (2.6a)
$=\bigl{(}-q^{2};q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}}$ (2.6b)
$=\bigl{(}q^{1/4};q^{1/2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{{k^{2}}/4}}{\bigl{(}q^{1/4};q^{1/4}\bigr{)}_{2k}},$ (2.6c)
$H(q)$ $=\bigl{(}-q^{2};q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+2)}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}}$ (2.6d)
$=\bigl{(}-q;q^{2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{k(k+1)}}{(-q;-q)_{2k+1}}$ (2.6e)
$=\bigl{(}q^{1/4};q^{1/2}\bigr{)}_{\infty}\sum_{k=0}^{\infty}\frac{q^{{k(k+2)}/4}}{\bigl{(}q^{1/4};q^{1/4}\bigr{)}_{2k+1}}.$ (2.6f)
@@
$A(q)=\sum_{k=0}^{\infty}\frac{q^{k^{2}}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}},\qquad B(q)=\sum_{k=0}^{\infty}\frac{q^{k(k+2)}}{\bigl{(}q^{4};q^{4}\bigr{)}_{k}}.$ (2.7)
Then, (2.6) can be written
$G\bigl{(}-q^{4}\bigr{)}=\bigl{(}q^{4};q^{8}\bigr{)}_{\infty}\frac{B(q)+B(-q)}{2},$ (2.8a)
$G(q)=\bigl{(}-q^{2};q^{2}\bigr{)}_{\infty}A(q),$ (2.8b)
$G\bigl{(}q^{16}\bigr{)}=\bigl{(}q^{4};q^{8}\bigr{)}_{\infty}\frac{A(q)+A(-q)}{2},$ (2.8c)
$H(q)=\bigl{(}-q^{2};q^{2}\bigr{)}_{\infty}B(q),$ (2.8d)
$H\bigl{(}-q^{4}\bigr{)}=\bigl{(}q^{4};q^{8}\bigr{)}_{\infty}\frac{A(q)-A(-q)}{2q},$ (2.8e)
$H\bigl{(}q^{16}\bigr{)}=\bigl{(}q^{4};q^{8}\bigr{)}_{\infty}\frac{B(q)-B(-q)}{2q^{3}}.$ (2.8f)
We may use (2.8a) and (2.8d) to eliminate $A$ and $B$ from the remaining equations. Writing also
$F(q)=\frac{\bigl{(}-q^{2};q^{2}\bigr{)}_{\infty}}{\bigl{(}q^{4};q^{8}\bigr{)}_{\infty}}=\frac{\bigl{(}q^{8};q^{8}\bigr{)}_{\infty}}{\bigl{(}q^{2};q^{2}\bigr{)}_{\infty}},$
we arrive at (1.2) or, equivalently, (1.3).
As was noted in the introduction, the identities (1.3) can be considered as recursions for computing the Taylor series of $G$ and $H$, starting from $G(0)=H(0)=1$; see (2.9) below. Hence, if we can verify that (1.2) also holds for the product sides
$\tilde{G}(q)=\frac{1}{\bigl{(}q,q^{4};q^{5}\bigr{)}_{\infty}},\qquad\tilde{H}(q)=\frac{1}{\bigl{(}q^{2},q^{3};q^{5}\bigr{)}_{\infty}},$
then the Rogers–Ramanujan identities $G=\tilde{G}$, $H=\tilde{H}$ follow. As Rogers observed, this reduces to classical theta function identities. We first write
$\tilde{G}(q)$ $=\frac{1}{\bigl{(}q,q^{4},q^{6},q^{9};q^{10}\bigr{)}_{\infty}}=\frac{\bigl{(}-q,-q^{9};q^{10}\bigr{)}_{\infty}}{\bigl{(}q^{4},q^{6};q^{10}\bigr{)}_{\infty}\bigl{(}q^{2},q^{18};q^{20}\bigr{)}_{\infty}}=\frac{\bigl{(}q^{8},q^{12};q^{20}\bigr{)}_{\infty}\theta\bigl{(}-q;q^{10}\bigr{)}}{(q^{2};q^{2})_{\infty}}$
@@
$\tilde{H}(q)=\frac{\bigl{(}q^{4},q^{16};q^{20}\bigr{)}_{\infty}\theta\bigl{(}-q^{3};q^{10}\bigr{)}}{\bigl{(}q^{2};q^{2}\bigr{)}_{\infty}}.$
@@
$\frac{\theta(z;q)+\theta(-z;q)}{2}=\sum_{k=-\infty}^{\infty}q^{k(2k-1)}z^{2k}=\theta\bigl{(}-qz^{2};q^{4}\bigr{)},$
$\frac{\theta(z;q)-\theta(-z;q)}{2}=-\sum_{k=-\infty}^{\infty}q^{k(2k+1)}z^{2k+1}=-z\theta\bigl{(}-q^{3}z^{2};q^{4}\bigr{)}.$
@@
$\frac{\tilde{G}(q)+\tilde{G}(-q)}{2}=\frac{\bigl{(}q^{8},q^{12};q^{20}\bigr{)}_{\infty}\bigl{(}-q^{12},-q^{28},q^{40};q^{40}\bigr{)}_{\infty}}{\bigl{(}q^{2};q^{2}\bigr{)}_{\infty}}.$
@@
$\bigl{(}-q^{12},-q^{28};q^{40}\bigr{)}_{\infty}=\frac{\bigl{(}q^{24},q^{56};q^{80}\bigr{)}_{\infty}}{\bigl{(}q^{12},q^{28};q^{40}\bigr{)}_{\infty}}=\frac{\bigl{(}q^{16},q^{24};q^{40}\bigr{)}_{\infty}}{\bigl{(}q^{12},q^{28};q^{40}\bigr{)}_{\infty}\bigl{(}q^{16},q^{64};q^{80}\bigr{)}},$
@@
$\frac{\tilde{G}(q)+\tilde{G}(-q)}{2}=\frac{\bigl{(}q^{8};q^{8}\bigr{)}_{\infty}}{\bigl{(}q^{2};q^{2}\bigr{)}_{\infty}\bigl{(}q^{16},q^{64};q^{80}\bigr{)}_{\infty}}=\frac{\bigl{(}q^{8};q^{8}\bigr{)}_{\infty}}{\bigl{(}q^{2};q^{2}\bigr{)}_{\infty}}\tilde{G}\bigl{(}q^{16}\bigr{)}.$
The remaining parts of (1.2), with $G$ and $H$ replaced by $\tilde{G}$ and $\tilde{H}$, follow similarly. Alternatively, the identities (1.3) can be obtained directly as special cases of
$\theta(z;q)=\theta\bigl{(}-qz^{2};q^{4}\bigr{)}-z\theta\bigl{(}-q^{3}z^{2};q^{4}\bigr{)}.$
@@
$F(q)=\sum_{k=0}^{\infty}f_{k}q^{2k},\qquad G(q)=\sum_{k=0}^{\infty}g_{k}q^{k},\qquad H(q)=\sum_{k=0}^{\infty}h_{k}q^{k}$
into (1.3) gives
$g_{2k}=\sum_{j=0}^{\lfloor k/8\rfloor}f_{k-8j}g_{j},\qquad$ $g_{2k+1}=\sum_{j=0}^{\lfloor k/2\rfloor}(-1)^{j}f_{k-2j}h_{j},$ (2.9a)
$h_{2k}=\sum_{j=0}^{\lfloor k/2\rfloor}(-1)^{j}f_{k-2j}g_{j},\qquad$ $h_{2k+1}=\sum_{j=0}^{\lfloor(k-1)/8\rfloor}f_{k-8j-1}g_{j}.$ (2.9b)
It is interesting to note that all the coefficients involved here have partition-theoretic interpretations. This is straight-forward for $f_{n}$, and very well-known for $g_{n}$ and $h_{n}$. This leads to the following result, which can be viewed as a combinatorial reformulation of (1.3).
@@
Let $f_{n}$ be the number of partitions of $n$ into parts not divisible by $4$. Let $g_{n}$ be the number of partitions of $n$ into parts congruent to ±$\pm 1$ $\operatorname{mod}$ $5$ or, equivalently, the number of partitions of $n$ without repeated or consecutive parts. Let $h_{n}$ be the number of partitions of $n$ into parts congruent to ±$\pm 2$ $\operatorname{mod}$ $5$ or, equivalently, the number of partitions of $n$ without repeated or consecutive parts and not containing $1$ as a part. Then, the recursions (2.9) hold.
3 Some additional results
Having proved the Rogers–Ramanujan identities, we can sum the six series in (2.6). This was of course clear to Rogers, and leads to the identities given in Slater’s list [12, 19, 20] as number 99, 20, 98, 16, 94 and 96, respectively (as is noted in [19], 98 is equivalent to 79, and 94 to 17).